# On the composable security of weak coin flipping

### v2

@@ B.3 Proof of Corollary 11
We denote the fractional part of $z$ as a binary string $b_{1}^{n}$. We denote the set $I=\{i\in\mathbb{N}\mid b_{i}=1\}$ as a string $a_{1}^{|I|}$ and the set $I^{\prime}=\{i\in\mathbb{N}\mid b_{i}=0\}$ as a string ${a^{\prime}}_{1}^{|I^{\prime}|}$ in ascending order. The $i$th WCF protocol $\pi_{i}$ as well as the other WCF protocols $\pi_{1}\|...\|\pi_{i-1}$ and $\pi_{i+1}\|...\|\pi_{n}$ output classical bits and follow strict temporal orders. Therefore, we obtain global security of $\pi_{i}$ according to Theorem 10
$\Pr\left[C_{i}=0|C_{1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (66)
$\Pr\left[C_{i}=1|C_{1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon.$ (67)
We first argue that a cheating party needs to maximize the winning probability of each $\pi_{i}$ in order to maximize the final winning probability. At each step $\pi_{i}$, if $b_{i}=1$, Alice wins with certainty if $C_{i}=0$ while wins with a probability smaller than $1$ if $C_{i}=1$; if $b_{i}=0$, Alice wins with a non-zero probability if $C_{i}=0$ while loses with certainty if $C_{i}=1$. Therefore, a cheating Alice strives to win at each step $\pi_{i}$ in order to finally win with the highest probability. It is the same for Bob. In conclusion, considering global security in Inequality (67), Alice’s winning probability is no larger than the case when
$\Pr\left[C_{i}=0|C_{1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon,$ (68)
and Bob’s winning probability is no larger than the case when
$\Pr\left[C_{i}=1|C_{1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon.$ (69)
@@
For a cheating Alice,
Pr[Alice $\Pr[\text{Alice wins}]$ $=\sum_{i=1}^{|I|}\Pr\left[C_{1}^{a_{i}}=(b_{1}^{a_{i}-1},0)\right]$ (70)
$=\sum_{i=1}^{|I|}\Pr\left[C_{a_{i}}=0|C_{1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\prod_{j=1}^{a_{i}-1}\Pr\left[C_{j}=b_{j}|C_{1}^{j-1}=b_{1}^{j-1}\right]$ (71)
@@
where $C_{1}^{0}=b_{1}^{0}$ is a certain event and Inequality (72) follows from Inequality (67). Similarly, for a cheating Bob, we obtain
Pr[Bob $\Pr[\text{Bob wins}]$ $=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{1}^{a^{\prime}_{i}}=(b_{1}^{a^{\prime}_{i}-1},1)\right]$ (73)
$=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{a^{\prime}_{i}}=1|C_{1}^{a^{\prime}_{i}-1}=b_{1}^{a^{\prime}_{i}-1}\right]\prod_{j=1}^{a^{\prime}_{i}-1}\Pr\left[C_{j}=b_{j}|C_{1}^{j-1}=b_{1}^{j-1}\right]$ (74)
@@
where we again apply Inequality (67) in (75). Then the bias is
$\epsilon^{\prime}_{A}$ $=|\Pr[\text{Alice wins}]-z|\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1}-\frac{1}{2^{a_{i}}}\leq\sum_{i=1}^{|I|}\frac{2(a_{i}-2i+2)}{2^{a_{i}}}\epsilon+o(\epsilon).$ (76)
We have omitted the absolute value because the winning probability of a cheating Alice is larger than that of an honest Alice, Pr[Alice $\Pr[\text{Alice wins}]\geq z$. Note that
@@
$\epsilon^{\prime}_{A}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+o(\epsilon)\leq\epsilon+o(\epsilon).$ (78)
@@
$\epsilon^{\prime}_{A}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{|I|}\frac{(a_{i}-2)}{2^{a_{i}-2}}\epsilon+o(\epsilon)\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{\infty}\frac{i-2}{2^{i-2}}\epsilon+o(\epsilon)\leq 2\epsilon+o(\epsilon).$ (79)
In conclusion, for Alice,
@@
Similarly for Bob,
$\epsilon^{\prime}_{B}=|\Pr[\text{Bob wins}]-z|\leq 2\epsilon+o(\epsilon).$ (81)
@@
$\epsilon^{\prime}=\max\{\epsilon^{\prime}_{A},\epsilon^{\prime}_{B}\}\leq 2\epsilon+o(\epsilon)$ (82)

### v3

@@ B.3 Proof of Corollary 11
Corollary.
The intuition for the proof of Corollary 11 is as follows. The probability of one party ultimately winning is a continuous function of the probabilities of winning in each round. Thus, a small bias in each round necessarily leads to a small overall bias.
Denote the fractional part of $z$ as a binary string $b_{1}^{n}$ (with finite precision), that is, $z=0.b_{1}b_{2}\dots b_{n}$. Define the sets $I\coloneqq\{i\in[n]\mid b_{i}=0\}$ and $I^{\prime}\coloneqq\{i\in[n]\mid b_{i}=1\}$. The elements of $I$ and $I^{\prime}$ form the sequences $(a_{1},\dots,a_{|I|})$ and $(a^{\prime}_{1},\dots,a^{\prime}_{|I^{\prime}|})$, respectively, arranged in ascending order. The $i$th WCF protocol $\pi_{i}$ as well as the other WCF protocols $\pi_{1}\|...\|\pi_{i-1}$ and $\pi_{i+1}\|...\|\pi_{n}$ output classical bits and follow strict temporal orders. Therefore, for $i=2,3,\dots,n$ and any bit sequence $c_{1}^{i-1}$, we have global security of $\pi_{i}$ according to Theorem 10:
$\Pr\left[C_{A,i}=1|C_{A,1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (78)
$\Pr\left[C_{B,i}=0|C_{B,1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (79)
where $C_{B(A),1}^{i}=C_{B(A),1}C_{B(A),2}\dots C_{B(A),i}$.
We first argue that a cheating party needs to maximize the winning probability of each game $\pi_{i}$ in order to maximize the final winning probability. At each step $\pi_{i}$, if $b_{i}=0$, Bob wins the final game with certainty if $C_{A,i}=1$ and wins the final game with a probability smaller than $1$ if $C_{A,i}=0$; if $b_{i}=1$, Bob wins with non-zero probability if $C_{A,i}=0$ while loses with certainty if $C_{A,i}=1$. Therefore, a cheating Bob strives to win at each round of $\pi_{i}$ in order to finally win with the highest probability. The similar argument holds for cheating Alice. In conclusion, Bob’s winning probability is no larger than the case when
$\forall i\in[n],\Pr\left[C_{A,i}=1|C_{A,1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon,$ (80)
and Alice’s winning probability is no larger than the case when
$\forall i\in[n],\Pr\left[C_{B,i}=0|C_{B,1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon.$ (81)
@@
According to the program of $\tau_{A}$ in Figure 7, Bob can win the final game only if the loop breaks at the $a$th round where $b_{a}=0$, or equivalently, $a\in I$. Then, his winning probability can be derived as
Pr[Bob $\Pr[\text{Bob wins}]$ $=\sum_{a\in I}\Pr[\text{Loop breaks at $a$th round and $C_{A,a}=1$}]$ (82)
$=\sum_{i=1}^{|I|}\Pr\left[C_{A,1}^{a_{i}}=(b_{1}^{a_{i}-1},1)\right]$ (83)
$=\sum_{i=1}^{|I|}\Pr\left[C_{A,a_{i}}=1|C_{A,1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\cdot\prod_{j=1}^{a_{i}-1}\Pr\left[C_{A,j}=b_{j}|C_{A,1}^{j-1}=b_{1}^{j-1}\right]$ (84)
$=\sum_{i=1}^{|I|}\Pr\left[C_{A,a_{i}}=1|C_{A,1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\cdot\prod_{j\in\{{a}_{k}\}_{k=1}^{i-1}}\Pr\left[C_{A,j}=0|C_{A,1}^{j-1}=b_{1}^{j-1}\right]\cdot$
$\hphantom{=}\prod_{j^{\prime}\in[a_{i}-1]\backslash\{{a}_{k}\}_{k=1}^{i-1}}\Pr\left[C_{A,j^{\prime}}=0|C_{A,1}^{j^{\prime}-1}=b_{1}^{j^{\prime}-1}\right]$ (85)
@@
where $C_{A,1}^{0}=b_{1}^{0}$ is defined as a certain event and Inequality (86) follows from Inequality (78) and (80). Similarly, for cheating Alice, we obtain
Pr[Alice $\Pr[\text{Alice wins}]$ $=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{B,1}^{a^{\prime}_{i}}=(b_{1}^{a^{\prime}_{i}-1},0)\right]$ (87)
$=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{B,a^{\prime}_{i}}=0|C_{B,i}^{a^{\prime}_{i}-1}=b_{1}^{a^{\prime}_{i}-1}\right]\prod_{j=1}^{a^{\prime}_{i}-1}\Pr\left[C_{B,j}=b_{j}|C_{B,1}^{j-1}=b_{1}^{j-1}\right]$ (88)
@@
Then the bias achieved by Bob is
$\epsilon^{\prime}_{B}$ $=|\Pr[\text{Bob wins}]-(1-z)|$ (90)
$\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1}-\frac{1}{2^{a_{i}}}$ (91)
$\leq\sum_{i=1}^{|I|}\frac{2(a_{i}-2i+2)}{2^{a_{i}}}\epsilon+o(\epsilon).$ (92)
We have omitted the absolute value because the winning probability of a cheating Bob is always larger than that of an honest Bob, Pr[Bob $\Pr[\text{Bob wins}]\geq 1-z$. Note that
@@
$\epsilon^{\prime}_{B}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+o(\epsilon)\leq\epsilon+o(\epsilon).$ (94)
@@
$\epsilon^{\prime}_{B}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{|I|}\frac{(a_{i}-2)}{2^{a_{i}-2}}\epsilon+o(\epsilon)\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{\infty}\frac{i-2}{2^{i-2}}\epsilon+o(\epsilon)\leq 2\epsilon+o(\epsilon).$ (95)
In conclusion, for cheating Bob,
$\epsilon^{\prime}_{B}=|\Pr[\text{Bob wins}]-(1-z)|\leq 2\epsilon+o(\epsilon).$ (96)
Similarly, for cheating Alice,
@@
$\epsilon^{\prime}=\max\{\epsilon^{\prime}_{A},\epsilon^{\prime}_{B}\}\leq 2\epsilon+o(\epsilon).$ (98)